The point

(π/6 , 1/2)


is on a curve and at any point on the curve, the slope of the tangent line is given by

−2 sin(2x)


. Find an equation of the curve.



1
Expert's answer
2021-05-27T17:27:49-0400

Given point is (π/6,1/2)


Given slope of tangent line is -2sin(2x)


So, dy/dx = -2sin(2x)


dy = -2sin(2x) dx


Integrating on both sides with respect to x


∫ dy = (-2) ∫ sin(2x) dx


y = (-2)(-1/2) cos(2x) + C


y = cos(2x) + C


Put (π/6,1/2) in above curve


(1/2) = cos(2*π/6) + C


(1/2) = (1/2) + C


C = 0


Thus, The equation of curve is y = cos(2x)